OpenStudy (anonymous):

what is the slope of a line perpendicular to 3x+2y=19? plz somebody help

OpenStudy (solomonzelman):

\(\normalsize\color{blue}{ 3x+2y=19}\) comert this to `y=mx+b` You will see the `m` in your equation, then take a negative multiplicative inverse of the `m` that you get.

OpenStudy (solomonzelman):

*convert, not comert

OpenStudy (anonymous):

can u do me a favour ? just solve it if u can ...

OpenStudy (trr54):

Or u could just do -A/B to find the slope in standard form

OpenStudy (anonymous):

if a line is perpendicular that means that it is totally opposite

OpenStudy (trr54):

It would be 2/3

OpenStudy (anonymous):

@Trr54 the ans is right but im having prob in solving it

OpenStudy (trr54):

B/c if u do -A/B it makes it into -3/2 After that step u have to change it to perpendicular... Do u know how to do that?

OpenStudy (anonymous):

nope

OpenStudy (trr54):

What you do is u have to flip the number and then put a negative in front of it... Did tht make sense or do u need more of an example?

OpenStudy (anonymous):

ahh i got it maybe .... :P can u just explan that B/c ??? and A/B ???

OpenStudy (trr54):

-A/B... Tht means tht... 3x+2y=19 Like in ur problem it would mean putting the number next to the x on top of the number next to the y with a negative sign in the front -A/B is just a formula u have to memorize it... It's to find the slope of any equation in standard form without having to convert it

OpenStudy (trr54):

When using this formula u have to be absolutely sure tht the number is in standard form correctly: Ax+By=C

OpenStudy (trr54):

Did u get it? Or would u like some further explanation?

OpenStudy (anonymous):

i somehow got it.. but theres one confusion that value of A will 3 and B 2 ??